Summary of the Meeting of the PS Management Board
held on 06.05.2002
Present : B.W. Allardyce,
S. Baird, J. Boillot, M. Chanel, G. Daems, J.P. Delahaye (Chairman), B. Frammery,
R. Garoby, G. Geschonke, J. Gruber, H. Haseroth, E. Kugler, S. Maury, K.D. Metzmacher,
S.L. Neboux (Secretary), J.P. Potier, J.P. Riunaud, K. Schindl
Excused : A. Blas (Staff Association),
* * * * *
1. Follow-up of previous meeting and
- PSMB 29.04.2002 - no follow-up.
- Restructuring of Accelerator Sector :
- Nominations : the Directorate has decided to proceed with
the restructuring of the Accelerator Sector to take effect on 1.1.2003
and division leaders for the two new divisions will be proposed to
Council in June: S. Myers for the new PS/SL Division and P. Lebrun
for the new LHC Division. S. Myers has been asked by the DG to undertake
the task of determining the new structure and staffing and to define
work packages for the different groups, with the redeployment of personnel
and work packages to and from other CERN activities. The focus of
resources on LHC is the main purpose for restructuring, with the aim
to redeploy in the order of 100 FTE/year (full-time equivalent). J.P.
Delahaye has accepted the proposal from the DG and S. Myers to act
as Deputy Division Leader, as he is in full agreement with the proposed
nomination of division leader and in particular to assure a smooth
transition with the best use possible of PS expertise and equal treatment
between PS and SL, with a fair distribution of responsibilities.
- PS/SL Management Meeting : Many areas still need to be discussed
and it is proposed to hold a three-day meeting from 27-29 May, attended
by the "old" and "new" management of PS and SL
to address these points.
- Removals : both S. Myers and J.P. Delahaye do not see the
need for major removals following the merging of PS and SL. Depending
on the different activities, the Group Leader concerned will decide
if certain removals are necessary and possible in his group.
- Tunnel : as the restructuring will involve more travel between
the two laboratories, it is expected that more people will need an
authorization to use the tunnel. Authorization will be at the discretion
of the Group Leader concerned. G. Daems pointed out the conditions
and restrictions (see Admin. Procedure "Tunnel"
and the DSU documents "Rules
for Use" and the information
note summarising them). An EDH
request for tunnel transit was introduced last week.
- Safety : G. Daems reported on recent safety issues :
- Power suppy of septum 57 : a serious incident occurred last
week when a repair was being carried out on this high voltage power
supply and a member of the Power Group received an electric shock,
which fortunately did not do too much harm, but could have been extremely
serious. The voltage supply to the septum was cut, but no one realized
that this old equipment had a second high voltage power supply and
in addition the obligatory connection with a grounding rod was not
made. A study of the possible modifications to improve the safety
of this equipment is being launched. J.P. Delahaye would like to
profit from this incident to remind everyone in the division that
safety instructions must be followed at all times. A meeting will
take place this afternoon between the G. Daems, DSO, and TIS.
- Fire alarm system : as a by-product of the water problem
last week in Linac 2 due to heavy rain (see below), the system covering
the two linacs and the power house was discovered to be out of order.
Renovation of this part of the system had been stopped because
of the budget cuts. However, following this breakdown, ST has now
found funds to renovate the system.
- Asbestos in LEAR/LEIR : an analysis of the air was made last
week and is slightly higher than the limit. An external specialist
has been contacted and more measurements will be made to see if the
area has to be decontaminated.
- Space for LHC project : the space manager for the Accelerator
Sector, W. Erdt, has addressed a message to group leaders regarding space
requirements for stocking LHC project equipment. PS has no particular
needs and B. Allardyce will reply on behalf of the division.
- Absence : J.P. Delahaye will be absent from 8-15.05.2002 and
will be replaced by K. Schindl.
- ISOLDE beam alignment : E. Kugler spoke about the responsibilities
for beam alignments. J.P. Riunaud confirmed that the EST/SU Group is responsible
for all beam alignments at CERN, in both the technical and accelerator
sectors, following the regrouping of this activity many years ago. To
help EST/SU with their planning, notification of required alignments should
be given as far in advance as possible.
- AD : in view of the importance of the LHC scrubbing run at the
SPS, it was decided to concentrate on the setting-up of the LHC beam in
PSB and PS from Thursday, 2.5.2002. This decision almost certainly results
in a delay of the setting-up of the AD proton production beam and a possible
delay in the AD physics start-up. S. Maury said that the physicists should
be informed of such a delay as soon as possible. J.P. Delahaye hoped that
it would be possible to make up for lost time and preferred to review
the situation next week before deciding on postponing the start of the
AD physics run.
- Septum 57 : the damage to the power supply means that beam momentum
for the East Hall is reduced from 24 GeV/c to 20
GeV/c. A firm will inspect the damaged transformer of the power supply
this afternoon to see if the coils can be repaired or whether they have
to be replaced. K. Metzmacher fears that redoing the coils may take anything
from 2 weeks to 2 months.
- Booster beam : R. Garoby evoked the serious misunderstanding
concerning the schedule of the various Booster beams required for the
PS start-up. As the booster started operation three weeks beforehand,
all these beams were expected to be ready for the PS start-up. The PSMB
decided that communication between the PS and PSB teams should be improved
and a detailed schedule for beams should be published and regularly updated
on the web so that the information is available to everyone at any time.
Action : S. Baird.
- Linac flooding : Linac 2 being part of the LHC injector chain,
ST should intervene to deal with this problem in a more permanent way.
There are two causes : i) a leaky roof and ii) a leaky heat exchanger
in the Linac air conditioning system. Action : S. Baird.
- Camera on the controlled access surveillance system : no piquet
service is available to deal with breakdowns. Repairs have to wait until
the next working day. An operator must go physically to the door to give
access if the camera is out of order.
3. Report on the IPHI Workshop
R. Garoby gave a summary
of the SPL
Mini-Workshop for the Definition of the Parameters of the RFQ and Chopper
Test Set-up, which was held at CERN from 25 to 26 April 2002 in the
context of the CEA-IN2P3-CERN collaboration on IPHI
(Injecteur de Protons de Haute Intensité). This collaboration will
be very beneficial in preparing the front-end of the planned 120 MeV injector
linac for the PSB and, at a later stage, the SPL.
It is planned to hold another meeting in Saclay in October 2002.
4. Report on the MCOG Meeting
H. Haseroth gave a presentation
on the Muon Concertation and Oversight Committee (MCOC) held at CERN in
April . A European Group has been set up - EMCOG (European Muon Concertation
and Oversight Group) - with C. Wyss as chairperson and A. Blondel as Scientific
Secretary. The next meeting will take place in London on 6-7 July 2002 between
NuFact'02 and MICE
(Muon Ionization Cooling Experiment). J.P. Delahaye emphasizes that the
manpower CERN can make available for NuFact
is extremely reduced (2.3 to 2.5 FTE/year), but that resources from other
laboratories would be very welcome.
| PS Division | PSMB
Meetings | Other PS Meetings |
7 May, 2002